OpenStudy (anonymous):

Express the equation in the logarithmic form of logC=D. (That is, the logarithm is base 10)

OpenStudy (anonymous):

\[10^{-5}=e-5\]

OpenStudy (anonymous):

So what are the values of C and D based on \[\log_{10}C=D \]

OpenStudy (anonymous):

\[\log_{10}(e-5) = -5 \]

OpenStudy (anonymous):

The value (e-5) is not the correct value of C.

OpenStudy (turingtest):

o-0 it should be...

OpenStudy (anonymous):

Common logs with an exponential? log10(10^-5)=log(e^-5) loga(a)=1 1(-5)-log(e^-5) -5=loge^-5

OpenStudy (anonymous):

I know. But the website says its wrong.

OpenStudy (turingtest):

is it really e-5 on the right?

OpenStudy (turingtest):

not e^(-5) as suggested above?

OpenStudy (anonymous):

The equation on the website is\[10^{-5}=1e-05\]

OpenStudy (anonymous):

well then that changes a bit...

OpenStudy (turingtest):

0.5 ?

OpenStudy (anonymous):

I don't know why there is a 0 in front of the 5

OpenStudy (anonymous):

the 1e - 05 is calculator notation for 10^(-5) [rather sloppy notation for it]

OpenStudy (turingtest):

Oh! lol

OpenStudy (anonymous):

so we have\[\log_{10}(10^{-5}) = -5 \]

OpenStudy (anonymous):

Ah that makes much more sense.

OpenStudy (anonymous):

Thanks a lot! I was so confused

OpenStudy (anonymous):

You know as I put down log10(e−5)=−5 to begin with I thought to myself, well that's not right as e-5 is negative and you can't have a log of a negative number.